# Clustering of DESI galaxies split by thermal Sunyaev-Zeldovich effect

### v1

@@ 2.1 Galaxy catalog: DESI DR1 LRG
The Dark Energy Spectroscopic Instrument (DESI DESI Collaboration et al., 2016b, 2022) is a robotic fiber spectrograph operating on the 4-meter Mayall telescope at Kitt Peak National Observatory. It conducts a 5-year galaxy survey of approximately a third of the sky, which will yield nearly 40 million galaxy and quasar redshifts (Levi et al., 2013; DESI Collaboration et al., 2016a). This efficiency is enabled by instrumentation and software innovations, including a focal plane design with the robotic positioners for approximately 5000 optical fibers (Silber et al., 2023; Poppett et al., 2024), an optical corrector widening the field of view to $\sim 3^{\circ}$, a comprehensive spectroscopic reduction pipeline (Guy et al., 2023) and a pipeline to adjust observation planning and optimization as survey progresses (Schlafly et al., 2023).
DESI’s scientific program has been successfully validated (DESI Collaboration et al., 2024b) alongside the Early Data Release (DESI Collaboration et al., 2024c). The key results based on Data Release 1 (DESI Collaboration et al., 2025c, recently made public) include clustering catalogs and two-point statistics measurements (DESI Collaboration et al., 2025g), baryon acoustic oscillation (BAO) distance measurements from galaxies, quasars (DESI Collaboration et al., 2025f) and Lyman-$\alpha$ forest (DESI Collaboration et al., 2025d) along with their detailed implications for cosmological models (DESI Collaboration et al., 2025e); and full-shape clustering of galaxies and quasars (DESI Collaboration et al., 2024a) with their cosmological analysis (DESI Collaboration et al., 2025h). Furthermore, updated BAO measurements from Lyman-$\alpha$ (DESI Collaboration et al., 2025b), galaxies and quasars with accompanying cosmological interpretations (DESI Collaboration et al., 2025a) are also available, based on Data Release 2 (DESI Collaboration, 2026).
We use Luminous Red Galaxies (LRG; Zhou et al., 2023a) from the DESI DR1 clustering catalog (DESI Collaboration et al., 2025g, c). We chose LRG because their redshift range ($0.4<z<1.1$) covers the peak in redshift distribution of Sunyaev-Zeldovich clusters (ACT/DES/HSC Collaboration et al., 2025). We discard galaxies at $z>0.85$ because then both the LRG density (DESI Collaboration et al., 2025g) and the SZ cluster abundance (ACT/DES/HSC Collaboration et al., 2025) drop.
@@ 2.2 Sunyaev-Zeldovich map: ACT DR6
We use the thermal Sunyaev-Zeldovich Compton $y$ parameter map from ACT Data Release 6 (Coulton et al., 2024), which builds upon Planck data (Planck Collaboration et al., 2016a). We rely on the accompanying noise simulations (Atkins et al., 2023) to estimate the signal-to-noise ratio in our analysis.
@@ 3.1 Methodology and challenges
We divide the LRGs into subsamples (“SNR bins”) according to the signal-to-noise ratio at their location in the ACT DR6 + Planck Compton-$y$ map (Coulton et al., 2024). First, we match the positions of the LRGs from the DESI clustering catalog to the pixels in the map [^1] The ACT DR6 map is provided as a rectangular grid in DEC and RA with a pixel side of 0.5 arcmin.. We only use the pixels with the apodized mask value above the threshold of 0.9. Second, we compute the pixel-level standard deviations in $y$ using the corresponding 304 Gaussian noise simulation (Atkins et al., 2023). Finally, we compute the signal-to-noise ratio in each pixel by dividing the $y$ value by its standard deviation.
@@ 3.2 Larger-scale clustering and galaxy bias
0 to 1 $\sigma$ subsample appears very similar to the full LRG sample. This is not explained by the vast majority of galaxies belonging to that bin, as 1 shows it contains just $\approx 1/3$ of our LRGs. Instead, the galaxies with negative signal and higher SNR compensate each other so that their environmental characteristics average close to this $[0,1)$ $\sigma$ bin. Interestingly, this similarity holds in all other aspects we consider in this paper.
@@ 4.1 Simple $y$ signal map from an $N$-body simulation
Using Eqs.˜7, 8 and 11, we can arrive at a simple expectation for Compton-$y$ parameter from a halo. First, we obtain $y\propto\sigma_{3d}^{2}\int\rho_{\rm dm}dl$. Then, we can estimate $\int\rho_{\rm dm}dl\propto M_{\rm halo}/R_{\rm halo}^{2}$. A simplistic definition of a halo is a spherical region with a fixed high density [^10] However, constant density does not hold in modern halo finders, including CompaSO (Hadzhiyska et al., 2022)., so $M_{\rm halo}\propto R_{\rm halo}^{3}$. Assuming virialization, $\sigma_{3d}^{2}\propto M_{\rm halo}/R_{\rm halo}$. Putting all together gives $y\propto M_{\rm halo}$, although halos can violate the underlying assumptions (and we see a steeper relation in Sections˜4.6 and 12).
@@ 4.4 Mock clustering
Currently, we do not focus on the small-scale line-of-sight clustering (analogs of Figs.˜5 and 6). One of the reasons is that we do not model fiber assignment in our cubic mocks for simplicity. On large scales, the fiber assignment incompleteness effects can be mitigated quite well with weights (Bianchi et al., 2025; DESI Collaboration et al., 2025g), but on small scales, it is a much more difficult problem.
@@ 4.6 Relating halo mass to thermal Sunyaev-Zeldovich observables
We present mock correlations between halo mass and several tSZ quantities [^19] We include these properties for all galaxies in the supplementary material, 10.5281/zenodo.16943122, and encourage searching for other interesting relations. in Fig.˜12. We count galaxies from a single mock galaxy catalog created with the fiducial HOD model, having their halo mass from the mock catalog and and obtaining tSZ quantities by matching them to our map pixels. In the upper panel of Fig.˜12 we see a power-law relation between the halo mass and the unfiltered noiseless Compton-$y$. The power-law looks slightly steeper than $y\propto M_{\rm halo}$ (estimated with big simplifications in Section˜4.1; the index may be similar to $\approx 1.08$ from Equation (4) in ACT/DES/HSC Collaboration et al. 2025). The trend smears up for smaller halo masses; this may be caused by such halos appearing close (within the point spread function) to slightly larger halos. The middle panel of Fig.˜12 shows a range of SNR values centered on zero for lower halo masses (as expected from Gaussian noise) and a noticeable increase for larger halo masses.
@@
Figure 12: Histograms displaying correlations between halo mass and tSZ quantities in the corresponding pixels in our mock map (assuming ACT DR6 + Planck characteristics following Coulton et al., 2024). Top: 2D histogram using the noiseless $y$ parameter without filtering. Middle: 2D histogram using the tSZ SNR with our fiducial filter and simplistic noise. Bottom: 1D histograms of halo mass in SNR (blue) or signal (orange) bins (the two distributions are very similar; both observables include our fiducial filter and simplistic noise). All histograms are weighted by galaxy counts from one mock catalog with the fiducial HOD model (see Section˜4.5).
A well-known or calibrated halo mass to tSZ observables relation can be used to make theoretical predictions about clustering, particularly on larger scales. Furthermore, if this relation is less dependent on cosmology than halo mass-density relations, it would enable more robust cosmological inference beyond the 2-point function. A theoretical or semi-analytical model would be a very attractive alternative to a simulation-based model (like for density-split clustering in Cuesta-Lazaro et al., 2024) requiring a large number of high-quality mocks.
@@ Acknowledgments
DESI Data Release 1 is available at https://data.desi.lbl.gov/doc/releases/dr1/. ACT Data Release 6 maps and noise simulations used in this work are available at https://portal.nersc.gov/project/act/dr6_nilc/. All the points from the figures, together with some additional data and the source code, are available at 10.5281/zenodo.16943122.

### v2

@@ 2.1 Galaxy catalog: DESI DR1 LRG
The Dark Energy Spectroscopic Instrument (DESI DESI Collaboration et al., 2016b, 2022) is a robotic fiber spectrograph operating on the 4-meter Mayall telescope at Kitt Peak National Observatory. Since 2021, it conducts an 8-year galaxy survey of approximately 40% of the sky, which will yield nearly 63 million galaxy and quasar redshifts (compared to the original forecast of 39 million for the 5-year program in Levi et al., 2013; DESI Collaboration et al., 2016a). This efficiency is enabled by instrumentation and software innovations, including a focal plane design with the robotic positioners for approximately 5000 optical fibers (Silber et al., 2023; Poppett et al., 2024), an optical corrector widening the field of view to $\sim 3^{\circ}$, a comprehensive spectroscopic reduction pipeline (Guy et al., 2023) and a pipeline to adjust observation planning and optimization as the survey progresses (Schlafly et al., 2023).
DESI’s scientific program has been successfully validated (DESI Collaboration et al., 2024b) alongside the Early Data Release (DESI Collaboration et al., 2024c). The key results based on Data Release 1 (DESI Collaboration et al., 2025c) include clustering catalogs and two-point statistics measurements (DESI Collaboration et al., 2025g), baryon acoustic oscillation (BAO) distance measurements from galaxies, quasars (DESI Collaboration et al., 2025f) and Lyman-$\alpha$ forest (DESI Collaboration et al., 2025d) along with their detailed implications for cosmological models (DESI Collaboration et al., 2025e); and full-shape clustering of galaxies and quasars (DESI Collaboration et al., 2024a) with their cosmological analysis (DESI Collaboration et al., 2025h). Furthermore, updated BAO measurements from Lyman-$\alpha$ (DESI Collaboration et al., 2025b), galaxies and quasars with accompanying cosmological interpretations (DESI Collaboration et al., 2025a) are also available, based on Data Release 2 (DESI Collaboration, 2026).
We use Luminous Red Galaxies (LRG; Zhou et al., 2023a) from the DESI DR1 clustering catalog covering 5,740 deg2 (DESI Collaboration et al., 2025g). We chose LRG because their redshift range ($0.4<z<1.1$) covers the peak in redshift distribution of Sunyaev-Zeldovich clusters (ACT/DES/HSC Collaboration et al., 2025). We discard galaxies at $z>0.85$ because then both the LRG density (DESI Collaboration et al., 2025g) and the SZ cluster abundance (ACT/DES/HSC Collaboration et al., 2025) drop.
@@ 2.2 Sunyaev-Zeldovich map: ACT DR6
We use the thermal Sunyaev-Zeldovich Compton $y$ parameter map from ACT Data Release 6 covering $\approx 13{,}000$ deg2 (Coulton et al., 2024), which builds upon Planck data (Planck Collaboration et al., 2016a). We rely on the accompanying noise simulations (Atkins et al., 2023) to estimate the signal-to-noise ratio in our analysis. We primarily work in the $\approx 3{,}510$ deg2 area of overlap between DESI DR1 LRG and ACT DR6 footprints.
@@ 3.1 Methodology and challenges
We divide the LRGs into subsamples (“SNR bins”) according to the signal-to-noise ratio at their location in the ACT DR6 + Planck Compton-$y$ map (Coulton et al., 2024). First, we match the positions of the LRGs from the DESI clustering catalog to the pixels in the map [^1] The ACT DR6 map is provided as a rectangular grid in DEC and RA with a pixel side of 0.5 arcmin.. We only use the pixels with the apodized mask value above the threshold of 0.9 (this slightly reduces the footprint from $\approx 3{,}510$ deg2 to $\approx 3{,}420$ deg2). Second, we compute the pixel-level standard deviations in $y$ using the corresponding 304 Gaussian noise simulation (Atkins et al., 2023). Finally, we compute the signal-to-noise ratio in each pixel by dividing the $y$ value by its standard deviation.
@@ 3.2 Larger-scale clustering and galaxy bias
The 0 to 1 $\sigma$ subsample appears very similar to the full LRG sample. This is not explained by the vast majority of galaxies belonging to that bin, as 1 shows it contains just $\approx 1/3$ of our LRGs. Instead, the galaxies with negative signal and higher SNR compensate each other so that their environmental characteristics average close to this $[0,1)$ $\sigma$ bin. Interestingly, this similarity holds in all other aspects we consider in this paper.
@@ 4.1 Simple $y$ signal map from an $N$-body simulation
Using Eqs.˜7, 8 and 11, we can arrive at a simple expectation for (central) Compton-$y$ parameter from a halo. First, we obtain $y\propto\sigma_{3d}^{2}\int\rho_{\rm dm}dl$. Then, we can estimate $\int\rho_{\rm dm}dl\propto M_{\rm halo}/R_{\rm halo}^{2}$. A simplistic definition of a halo is a spherical region with a fixed high density [^10] However, constant density does not hold in modern halo finders, including CompaSO (Hadzhiyska et al., 2022)., so $M_{\rm halo}\propto R_{\rm halo}^{3}$. Assuming virialization, $\sigma_{3d}^{2}\propto M_{\rm halo}/R_{\rm halo}$. Putting all together gives $y\propto M_{\rm halo}$, although halos can violate the underlying assumptions (and we see a steeper relation in Sections˜4.6 and 12).
@@
The naive expectation $y\propto M_{\rm halo}$ for the central Compton-$y$ parameter from a halo (using Eqs.˜7, 8 and 11) can hold for halos larger than the pixel, beam and filter size. For small, unresolved halos, there is an additional factor of halo area relative to the fixed pixel, beam or pixel area $\propto R_{\rm halo}^{2}\propto M_{\rm halo}^{2/3}$. Alternatively, Eq.˜16 can be used together with Eqs.˜8 and 11 again. Either way, we obtain a steeper scaling $y_{\rm map}\propto M_{\rm halo}^{5/3}$ as a simple expectation for smaller, unresolved halos.
@@ 4.2 Simple noise model
This treatment allows us to mimic the zero-mean Gaussian fluctuations caused by instrumental noise and some component-separation systematics. However, it does not capture systematic biases, deviations from normal distribution and potential correlations between the signal and the noise. As a result, negative SNR in our mocks only results from Gaussian smearing, whereas in data, systematic biases may also contribute.
@@ 4.4 Mock clustering
One of the real data systematics not modeled in our cubic mocks is fiber assignment incompleteness (Bianchi et al., 2025). As this effect should lead to the underrepresentation of denser regions, we might expect lower relative bias for positive tSZ SNR in data compared to mocks. However, in Fig.˜10, we see the opposite between 1 and $4\sigma$. This may be partially explained by the successful mitigation of fiber assignment incompleteness effects in large-scale clustering with weights (Bianchi et al., 2025; DESI Collaboration et al., 2025g) and on-sky separation ($\theta$) cut (Pinon et al., 2025), as we use both on the data side. Correcting for fiber assignment on small scales is a much more difficult problem, and this is a significant reason why we currently do not focus on the line-of-sight clustering (analogs of Figs.˜5 and 6).
@@ 4.5 Dependency on galaxy-halo connection
We keep the same tSZ map with varying HOD parameters, so the same halos remain in each SNR bin. $M_{\rm cut}$ and $\sigma$ change the number of LRG centrals and satellites in these halos. $M_{1}$, $\alpha$ and $\kappa$ only affect the number of satellites. $\alpha_{c}$ and $\alpha_{s}$ only alter redshift-space positions of centrals and satellites respectively. All these affect the redshift-space clustering of LRGs.
@@
The relative bias shifts change sign between the $0,1)~\sigma$ bin and the $1,2)~\sigma$ bin for all parameters except $\alpha_{c}$ [^20] For $\alpha_{c}$, the sign flips between $1,0)~\sigma$ and $0,1)~\sigma$, but all the differences have very low significance. in the bottom panel of Fig.˜11. The direction flip in itself is only natural, because the average relative bias (weighted by the number [^21] For data, it should be weighted by the sum of clustering weights. But in our cubic mocks, all the galaxies have weights set to 1. of galaxies in each corresponding SNR bin) should be 1222222As it is defined as the scaling of the full-sample auto-correlation function to best match the SNR bin-full sample cross-correlation functions. However, it is not clear to us why the change of sign is around $1\sigma$ tSZ detection significance specifically for most of the HOD parameters.
@@ 4.6 Relating halo mass to thermal Sunyaev-Zeldovich observables
Figure 12: Histograms displaying correlations between LRG host halo masses and tSZ quantities in the corresponding pixels in our mock map (assuming ACT DR6 + Planck characteristics following Coulton et al., 2024; Atkins et al., 2023). Top: 2D histogram using the noiseless $y$ parameter without filtering. Middle: 2D histogram using the tSZ SNR with our fiducial filter and simplistic noise. Bottom: 1D histograms of halo mass in SNR (blue) or signal (orange) bins (the two distributions are very similar; both observables include our fiducial filter and simplistic noise). All histograms are weighted by galaxy counts from one mock catalog with the fiducial HOD model (see Section˜4.5).
We present correlations between host halo mass and several tSZ quantities for mock galaxies [^23] We include these properties for all galaxies in the supplementary material, 10.5281/zenodo.16943122, and encourage searching for other interesting relations. in Fig.˜12. We count LRGs from a single mock galaxy catalog created with the fiducial HOD model, having their host halo mass listed in the mock catalog [^24] Therefore, the absence of $M_{\rm halo}\lesssim 10^{12.3}~M_{\odot}/h$ is not a limitation of halos in the AbacusSummit simulation but a reflection of the best-fit HOD model from Yuan et al. (2024). We tried plotting halo mass distributions, but found it much less informative due to a very strong contribution of smaller halos down to $\sim 10^{11}~M_{\odot}/h$ that would almost never host LRGs. The absence of $M_{\rm halo}\lesssim 10^{11}~M_{\odot}/h$ halos in AbacusSummit ”base” boxes we used might lead to misestimation of the mock tSZ signal, but their contribution is likely not very significant (as we naively expect $y_{\rm map}\propto M_{\rm halo}^{5/3}$ for each very low-mass halo from Section 4.1, and the more realistic power index may be even steeper)., and obtain tSZ quantities by matching them to our map pixels. In the upper panel of Fig.˜12 we see a power-law relation between the halo mass and the unfiltered noiseless Compton-$y$. The power-law looks slightly steeper than $y\propto M_{\rm halo}$ (estimated with big simplifications in Section˜4.1; the index may be similar to $\approx 1.08$ from Eq. (4) in ACT/DES/HSC Collaboration et al. 2025). The trend smears up for smaller halo masses; this may be caused by such halos appearing close (within the point spread function) to slightly larger halos. The middle panel of Fig.˜12 shows a range of SNR values centered on zero for lower halo masses (as expected from Gaussian noise and negligible signal) and a noticeable increase for larger halo masses.
@@
A well-known or calibrated relation between the galaxy host halo masses and tSZ observables can be used to make theoretical predictions about tSZ-split galaxy clustering, especially on larger scales. Furthermore, if this relation is less dependent on cosmology than halo mass-density relations, it would enable more robust cosmological inference beyond the 2-point function. A theoretical or semi-analytical model would be a very attractive alternative to a simulation-based model (like for density-split clustering in Cuesta-Lazaro et al., 2024) requiring a large number of high-quality mocks.
@@ Acknowledgments
DESI Data Release 1 is available at https://data.desi.lbl.gov/doc/releases/dr1/; the clustering catalogs used in this work can be found at https://data.desi.lbl.gov/public/dr1/survey/catalogs/dr1/LSS/iron/LSScats/v1.5/. ACT Data Release 6 maps and noise simulations used in this work are available at https://portal.nersc.gov/project/act/dr6_nilc/. All the points from the figures, together with some additional data and the source code, are available at 10.5281/zenodo.16943122.